OpenStudy (anonymous):

What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

The line has as standard form: y = ax + b. Find a and b Write that this line passes at point (-4, -4) and (4, 8) -4 = -4a + b 8 = 4a + b Add up the 2 equations 2b = 4 -> b = 2. y-coordinate of y-intercept is b = 2.